OpenStudy (anonymous):

A 75.0 mL volume of 0.200 M NH3 (Kb=1.8*10^-5) is titrated with 0.500 M HNO3. Calculate the pH after the addition of 25.0 mL of HNO3.

OpenStudy (anonymous):

anybody..please help me..urgent !

OpenStudy (anonymous):

try..

OpenStudy (rogue):

\[NH_3 + HNO_3 \rightarrow NH_4 NO_3\]\[n = MV\]Moles of NH3:\[n = 0.075 L * 0.200 M = 0.015 mol NH_3\]Moles of HNO3:\[n = 0.025 L * 0.500 M = 0.0125 mol HNO_3\]

OpenStudy (rogue):

HNO3 is a strong acid, so it will dissociate completely into protons & nitrate ions. The 0.0125 mol HNO3 will neutralize 0.0125 mol NH3 and form ammonium nitrate. 0.015 mol - 0.0125 mol = 0.0025 mol NH3 remaining

OpenStudy (rogue):

\[M = \frac {n}{V} = \frac {0.0025 mol NH_3}{0.100 L} = 0.025 Molar NH_3\]Now we can make an ice chart for this remaining amount of weak base.\[NH_3 + H_2O \rightarrow NH _{4}^{+} + OH^{-}\][Initial] 0.025 M________ 0 M_______ 0M [Change] -x +x +x [Final] 0.025 M - x x x

OpenStudy (rogue):

\[k_b = \frac {[OH^{-}][NH4_{4}^{+}]}{[NH3]}\]\[1.8 \times 10^{-5} = \frac {x^2}{0.025 M - x}\]\[x \approx 6.62 \times 10^{-4} M\]

OpenStudy (rogue):

x is the concentration of hydroxide ions, so we can get the pOH from that.\[pOH = - \log [OH]\]\[pOH = - \log_{10} 6.62 \times 10^{-4} \approx 3.18 \]\[pOH + pH = 14\]\[3.18 + pH = 14\]\[pH = 10.82\]

OpenStudy (rogue):

That was long! Hopefully I didn't make any mistakes because there were a lot of steps...